OpenStudy (anonymous):

Compute the Laplace Transform. y(0)=1 ; y'(0)=-1 y'' + 5y' + 3y = 0

OpenStudy (anonymous):

I evaluated it to L[y] = -(s+4) / (s^2 + 5s + 3) I am pretty certain my algebra is correct, I am just not sure how I would evaluate this further.

OpenStudy (freckles):

maybe I did some algebra wrong but I got \[Y(s)=\frac{s+4}{s^2+5s+3}\] anyways try completing the square on bottom \[Y(s)=\frac{s+4}{s^2+5s+(\frac{5}{2})^2+3-(\frac{5}{2})^2} \\ Y(s)=\frac{s+4}{(s+\frac{5}{2})^2-\frac{13}{4}}\]

OpenStudy (freckles):

there is a couple of other tricks you could use 5/2+3/2=4 and then you will need have to do a fancy multiplication of 1 not incredibly fancy, just a smidgen fancy

OpenStudy (anonymous):

ah got it, thank you very much!